OpenStudy (anonymous):

identify the inside function u= g(x) and the outside function y= f(u) y= f(g(x)), u = g(x) y = f(u) y = (6x-5)^4

OpenStudy (anonymous):

inside function is \[u=g(x)=6x-5\] outside function is \[f(u)=u^4\]

OpenStudy (anonymous):

so \[f(g(x)=f(6x-5) = (6x-5)^4\]

OpenStudy (anonymous):

oh ok cool i was wondering thanks i now how to do it now i just wasnt use to the terms

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

thats not all you have to solve for that question right i have to use the chain rule or no?